# Well behaved class of Heintzmann’s solution within $f(R,\,T)$ framework

### v1

@@ I Introduction
In this connection, it should be mentioned that scientists are stumped by the idea that the universe is growing faster than expected because of some sort of secret energy that has existed since the Big Bang. There may be a link between large-scale alteration of gravity and this late-time accelerated expansion of the cosmos Riess et al. (1998); Koyama (2016). Both the acceleration of expansion and the modification of gravity compared to General Relativity (GR), are caused actually by the presence of dark energy (DE), dark matter(DM), or a new form of matter, and interestingly these two scenarios are not in any way related, leading to numerous stellar models Copeland et al. (2006); Joyce et al. (2015). Dark energy can be best understood in terms of the cosmological constant, which satisfies cosmic data and represents a constant energy density of the vacuum. An in-depth search has been conducted for alternative theories of gravity that extend Einstein’s theory of gravity, due to the problematic nature of the cosmological constant Bhar and Rej (2023). Several researchers have suggested various modified theories of gravity, such as $f(R)$-gravity, $f(Q,T)$-gravity, $f(\mathcal{G},T)$-gravity, $f(R,\mathcal{G})$-gravity, unimodular gravity, teleparallel gravity, etc. Nojiri et al. (2016); García-Aspeitia et al. (2019); Nojiri and Odintsov (2007, 2011); Xu et al. (2019); Arora et al. (2020); Bahamonde et al. (2023); Atazadeh and Darabi (2014); De Felice et al. (2011). where $R$ indicates the Ricci scalar, $T$ denotes the trace of the stress-energy tensor and $\mathcal{G}$ indicates the Gauss-Bonnet invariant. T. Harko and his associates proposed $f(R,T)$ gravity theory in 2011 Harko et al. (2011). A generalization of the gravitational Lagrangian of the $f(R)$ gravity serves as its foundation. An arbitrary function of the Ricci scalar $R$ and the trace of the energy-momentum tensor $T$ determine the gravitational Lagrangian of the standard Hilbert-Einstein action in the $f(R,T)$ theory of gravity.
@@
Consequently, the motion of the test particle deviates from the geodesic path because there is an additional force perpendicular to the four-velocity directions. In this updated $f(R,T)$ hypothesis, the cosmic acceleration arises from both the matter content and the geometrical aspect Chakraborty (2013). The shift from a matter-dominated era to an accelerated phase is explored by Houndjo HOUNDJO (2012), who picked $f(R,T)$ as $f_{1}(R)+f_{2}(T)$, where $f_{1}(R)$ is a function of the Ricci scalar and the $f_{2}(T)$ can be considered as a matter correction term to $f(R)$ gravity. The trace of energy-momentum tensor disappears for ultrarelativistic fluids; consequently, these constituents of matter do not affect the function of $f(R,T)$. To address this shortcoming, a generalization of this theory has also been developed, incorporating a new invariant, or $R_{\alpha\beta}T_{\alpha\beta}$ Sharif and Zubair (2013); Odintsov and Sáez-Gómez (2013); Shabani and Farhoudi (2014). Functions of $f(R,T)$ that are distinguishable either minimally or non-minimally, into arbitrary functions of the trace of the energy-momentum tensor, $h(T)$, and an arbitrary function of the Ricci scalar, $g(R)$, like $g(R)+h(T)$, $g(R)h(T)$, $g(R)\{1+h(T)\}$, etc. The cosmological solution to the second type of assumption ($g(R)h(T)$) cannot be consistent.
@@
Numerous fields have already seen impressive outcomes from the $f(R,T)$ gravity such as massive pulsars dos Santos Jr et al. (2019); Moraes et al. (2016), dark energy Sun and Huang (2016), dark matter Zaregonbadi et al. (2016), wormholes Moraes and Sahoo (2019); Moraes et al. (2019); Elizalde and Khurshudyan (2018); Moraes and Sahoo (2018); Sahoo et al. (2020a), gravitational waves Sharif and Siddiqa (2019); Alves et al. (2016), bouncing cosmology Bhattacharjee and Sahoo (2020); Sahoo et al. (2020b) etc. In their discussion of the non-static spherically symmetric line element, Sharif et al. addressed the stability of a collapsing spherical body of an isotropic fluid distribution Sharif and Yousaf (2014). However, to solve the instability problem, Noureen et al. created a perturbation scheme for determining the collapse equation and a constraint on the adiabatic index for the Newtonian and post-Newtonian periods Noureen and Zubair (2015a) and additionally, the range of instability for an anisotropic background confined by zero expansion under the $f(R,T)$ theory has been developed by them Noureen and Zubair (2015b). Using the analytic solution of the Krori and Barua metric to the spherically symmetric anisotropic star, Zubair et al. examined the prospect of the development of compact stars in the $f(R,T)$ theory of gravity Zubair et al. (2016). The spherical equilibrium configuration of strange and polytropic stars under the $f(R,T)$ theory of gravity has been studied by Moraes et al Moraes et al. (2016). Ahmed et al. examined how gravitational lensing is affected by $f(R,T)$ gravity and contrasted their findings with the standard outcome of GR Alhamzawi and Alhamzawi (2016).
@@
In this work, we have considered the Heintzmann IIa ansatz for the particular $f(R,T)=R+2\chi T$ framework, where $\chi$ is some constant to explain a static, spherically symmetric arrangement connected to an isotropic fluid substance. H. Heintzmann Heintzmann (1969); Delgaty and Lake (1998) derived a method to derive new exact EFE solutions for an ideal isotropic fluid from the solutions obtained previously. Later, researchers discovered that it could be useful in a variety of astrophysical experiments. Heintzmann’s solution is so simple, making it easy to study general relativistic stars. Pant et al. Pant et al. (2010) introduced a well-behaved class of charged analogs of Heintzmann’s relativistic solutions. Thirukkanesh et al. Thirukkanesh et al. (2018) presented a generalized algorithm for EFE solutions that can be reduced to the Heintzmann IIa solution. Anisotropy and charge were added to the Heintzmann solution by Singh and Pant Singh and Pant (2016) to simulate dense compact objects. By using the gravitational decoupling technique, Estrada and Tello-Ortiz Estrada and Tello-Ortiz (2018) were able to obtain the new anisotropic solution by using this solution as an interior solution. Morales and Tello-Ortiz Morales and Tello-Ortiz (2018) developed charged anisotropic solutions to the EFEs by extending the Heintzmann solution by utilizing gravitational decoupling. Zubair et al. Zubair et al. (2021) extended the well-known charged isotropic Heintzmann solution to its anisotropic domain by using the gravitational decoupling approach. This Heintzmann solution has been taken into consideration in several other research projects Sharif and Majid (2020); Sharif and Waseem (2020); Andrade (2022) (and further references therein).
@@
This paper is organized as follows: In Section II, $f(R,T)$ gravity is defined mathematically in terms of isotropic matter distributions. A conceivable $f(R,\,T)$ gravity model is displayed in Section III. Section IV deals with boundary conditions; right here, we compare the interior and Schwarzschild’s exterior metrics to determine the values of the unknown constant for the selected values of our model parameters. Section V executes graphical analysis to investigate various physical properties and viability of compact stars. The last section is devoted to the conclusions. Throughout the investigation, we will mostly utilize a negative metric signature (+,,,)(+,-,-,-)( + , - , - , - ), and we will select natural units as $c=1=G$.
@@ II Interior Spacetime and Basic field Equations
We now employ a typical outcome, which is
@@
and the variation of the above action(1) can be written as:
@@
$f_{R}R_{\mu\nu}-\frac{1}{2}f(R,T)g_{\mu\nu}+(g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu})f_{R}$ $=$ $8\pi T_{\mu\nu}-f_{T}T_{\mu\nu}$ (6)
$-f_{T}\Theta_{\mu\nu}.$
@@
$\nabla^{\mu}T_{\mu\nu}$ $=$ $\frac{f_{T}}{8\pi-f_{T}}\left[(T_{\mu\nu}+\Theta_{\mu\nu})\nabla^{\mu}\ln f_{T}+\nabla^{\mu}\Theta_{\mu\nu}\right].$ (7)
From eqn.(7), we can check that $\nabla^{\mu}T_{\mu\nu}\neq 0$ if $f_{T}(R,T)\neq 0.$ Shortly thereafter, equation (7) demonstrates that the energy-momentum tensor is not conserved for the $f(R,\,T)$ theory of gravity, in contrast to general relativity. Moreover, when $f(R,T)=f(R)$, from eqn. (6) We derive the $f(R)$ gravity field equations.
@@
where $p$ denotes isotropic pressure under modified gravity and $\rho$ denotes matter density. The fluid velocity, $u^{\mu}$, fulfils the following equations: $u^{\mu}u_{\mu}=1$ and $u^{\mu}\nabla_{\nu}u_{\mu}=0$. As proposed by Harko et al. Harko et al. (2011), we have taken the matter Lagrangian that can be taken as $\mathcal{L}_{m}=-p$ and the expression of $\Theta_{\mu\nu}=-2T_{\mu\nu}-pg_{\mu\nu}.$
@@
$G^{0}_{0}:8\pi\rho^{\text{eff}}$ $=$ $\frac{\lambda^{\prime}}{r}e^{-\lambda}+\frac{1}{r^{2}}(1-e^{-\lambda}),$ (12)
$G^{1}_{1}:8\pi p^{\text{eff}}$ $=$ $\frac{1}{r^{2}}(e^{-\lambda}-1)+\frac{\nu^{\prime}}{r}e^{-\lambda},$ (13)
$G^{2}_{2}=G^{3}_{3}:8\pi p^{\text{eff}}$ $=$ $\frac{1}{4}e^{-\lambda}\left[2\nu^{\prime\prime}+\nu^{\prime 2}-\lambda^{\prime}\nu^{\prime}+\frac{2}{r}(\nu^{\prime}-\lambda^{\prime})\right].$ (14)
Here $\rho^{\text{eff}}$ and $p^{\text{eff}}$ are respectively the density and pressure in Einstein Gravity and
$\rho^{\text{eff}}$ $=$ $\rho+\frac{\chi}{8\pi}(3\rho-p),$ (15)
$p^{\text{eff}}$ $=$ $p-\frac{\chi}{8\pi}(\rho-3p),$ (16)
the prime indicates differentiation with respect to ‘r’. Using Eqs. (12)-(14), we get,
@@ III Exact Solution of our proposed Model for isotropic Stars
Utilizing the following metric potentials, one can compute the effective density and pressure as:
$\rho^{\text{eff}}$ $=$ $\frac{3B\bigg{\{}(3+Br^{2})(1+4Br^{2})^{3/2}+C(3+9Br^{2})\bigg{\}}}{16\pi(1+Br^{2})^{2}(1+4Br^{2})^{3/2}},$ (20)
$p^{\text{eff}}$ $=$ $-\frac{3B\bigg{\{}C+7BCr^{2}+3(-1+Br^{2})\sqrt{1+4Br^{2}}\bigg{\}}}{16\pi(1+Br^{2})^{2}\sqrt{1+4Br^{2}}},$ (21)
The physical parameters, like as effective pressure and density, must be obtained before so that we can investigate the full structure of stellar models.
By utilizing the $p^{\text{eff}}$ and $\rho^{\text{eff}}$ expressions from equations (15) and (16), we can derive the matter density and pressure expressions $\rho,\,p$ in modified gravity as,
@@ IV Exterior line element and matching conditions
[…]
$C$ $=$ $\frac{(1+4B\mathfrak{R}^{2})^{3/2}\Big{(}3\chi+6\pi-2B(\chi+3\pi)\mathfrak{R}^{2}\Big{)}}{2\pi+2B(3\chi+11\pi)\mathfrak{R}^{2}+7B^{2}(3\chi+8\pi)\mathfrak{R}^{4}}$
By selecting a preferred value of $B$ ($=0.001~{}\rm{km}^{-2}$), the constant values of $A$ and $C$, which are listed in Table 1 for various values of $\chi$, are determined using the approximate mass and radius of the compact star Vela X-1. Interestingly, the table shows that the value of $A$ remains unchanged, meaning that it is independent of $\chi$ throughout our model. On the contrary, the values of $C$ increase as $\chi$ increases.

### v2

@@ I Introduction
[…]
@@
Consequently, the motion of the test particle deviates from the geodesic path because there is an additional force perpendicular to the four-velocity directions. In this updated $f(R,T)$ hypothesis, the cosmic acceleration arises from both the matter content and the geometrical aspect Chakraborty (2013). The shift from a matter-dominated era to an accelerated phase is explored by Houndjo Houndjo (2012), who picked $f(R,T)$ as $f_{1}(R)+f_{2}(T)$, where $f_{1}(R)$ is a function of the Ricci scalar and the $f_{2}(T)$ can be considered as a matter correction term to $f(R)$ gravity. The trace of energy-momentum tensor disappears for ultrarelativistic fluids; consequently, these constituents of matter do not affect the function of $f(R,T)$. To address this shortcoming, a generalization of this theory has also been developed, incorporating a new invariant, or $R_{\alpha\beta}T^{\alpha\beta}$ Sharif and Zubair (2013); Odintsov and Sáez-Gómez (2013); Shabani and Farhoudi (2014a). Functions of $f(R,T)$ that are distinguishable either minimally or non-minimally, into arbitrary functions of the trace of the energy-momentum tensor, $h(T)$, and an arbitrary function of the Ricci scalar, $g(R)$, like $g(R)+h(T)$, $g(R)h(T)$, $g(R)\{1+h(T)\}$, etc. Later found that, the cosmological solution to the second type of pure non-minimal assumption $g(R)h(T)$ is not physically justified Shabani and Farhoudi (2013, 2014b).
@@
Several fields have already witnessed impressive outcomes from the $f(R,T)$ gravity. Such as application of $f(R,T)$ gravity for massive pulsars by Santos Jr. et al dos Santos Jr et al. (2019), Moraes et al Moraes et al. (2016); cosmological consequences without dark energy by Sun and Huang Sun and Huang (2016); dark matter theories by Zaregonbadi et al Zaregonbadi et al. (2016); investigations on wormholes by Moraes and his co-workers Moraes and Sahoo (2019); Moraes et al. (2019); Moraes and Sahoo (2018), Elizalde and Khurshudyan Elizalde and Khurshudyan (2018), Sahoo et al Sahoo et al. (2020a); study on gravitational waves by Sharif and Siddiqa Sharif and Siddiqa (2019), Alves et al Alves et al. (2016); bouncing cosmological scenario by Sahoo et al Sahoo et al. (2020b), Bhattacharjee and Sahoo Bhattacharjee and Sahoo (2020); and so on. In their discussion of the non-static spherically symmetric line element, Sharif et al. addressed the stability of a collapsing spherical body of an isotropic fluid distribution Sharif and Yousaf (2014). However, to solve the instability problem, Noureen et al. created a perturbation scheme for determining the collapse equation and a constraint on the adiabatic index for the Newtonian and post-Newtonian periods Noureen and Zubair (2015a) and additionally, the range of instability for an anisotropic background confined by zero expansion under the $f(R,T)$ theory has been developed by them Noureen and Zubair (2015b). Using the analytic solution of the Krori and Barua metric to the spherically symmetric anisotropic star, Zubair et al. examined the prospect of the development of compact stars in the $f(R,T)$ theory of gravity Zubair et al. (2016). The spherical equilibrium configuration of strange and polytropic stars under the $f(R,T)$ theory of gravity has been studied by Moraes et al Moraes et al. (2016). Ahmed et al. examined how gravitational lensing is affected by $f(R,T)$ gravity and contrasted their findings with the standard outcome of GR Alhamzawi and Alhamzawi (2016).
@@
In this work, we have considered the well-known Heintzmann IIa ansatz for the particular $f(R,T)=R+2\chi T$ framework, where $\chi$ is some arbitrary constant to explain a static, spherically symmetric arrangement connected to an isotropic fluid substance that represents compact spherical objects. In 1969, H. Heintzmann Heintzmann (1969); Delgaty and Lake (1998) developed a method to derive new exact EFE solutions for an ideal isotropic fluid from the solutions obtained previously. Later, researchers discovered that it could be useful in a variety of astrophysical experiments Andrade (2022a); Pradhan and Pant (2015). Several approaches have already been used to expand this solution and produce physically intriguing results. Heintzmann’s solution is so simple, making it easy to study general relativistic stars. To create new static and spherical solutions for isotropic fluid distributions, it acts as an initial solution. This is essential to comprehend the interior of relativistic compact objects, which might display variations from isotropy as a result of phase transitions, high densities, or strong magnetic fields. Moreover, The framework of gravitational decoupling, a technique to expand the solution space of Einstein’s field equations by taking into account more sources of gravity, makes use of the Heintzmann IIa solution and so it facilitates the investigation of more complicated compact object models. The framework of the Heintzmann IIa solution includes both the trace of the energy-momentum tensor and the Ricci scalar and so, compared to models that might only take one of these aspects into account, this dual inclusion enables a more thorough understanding of gravitational interactions. Assuming isotropic pressure inside the fluid sphere is the foundation of the Heintzmann IIa solution. Anisotropic pressure can, however, actually exist inside compact objects for a variety of reasons, including rotation, phase transitions, and strong magnetic fields. So, this could reduce how accurately the solution reflects actual objects. Overall, the Heintzmann IIa solution provides an innovative viewpoint on the stability and internal structure of compact objects in the universe and is a strong model inside the $f(R,T)$ framework. Pant et al. Pant et al. (2010) introduced a well-behaved class of charged analogs of Heintzmann’s relativistic solutions. Thirukkanesh et al. Thirukkanesh et al. (2018) presented a generalized algorithm for EFE solutions that can be reduced to the Heintzmann IIa solution. Anisotropy and charge were added to the Heintzmann solution by Singh and Pant Singh and Pant (2016) to simulate dense compact objects. By using the gravitational decoupling technique, Estrada and Tello-Ortiz Estrada and Tello-Ortiz (2018) were able to obtain the new anisotropic solution by using this solution as an interior solution. Morales and Tello-Ortiz Morales and Tello-Ortiz (2018) developed charged anisotropic solutions to the EFEs by extending the Heintzmann solution by utilizing gravitational decoupling. Zubair et al. Zubair et al. (2021) extended the well-known charged isotropic Heintzmann solution to its anisotropic domain by using the gravitational decoupling approach. This Heintzmann solution has been taken into consideration in several other research projects Sharif and Majid (2020); Sharif and Waseem (2020); Andrade (2022b) (and further references therein).
For this investigation, we have chosen a particular compact star candidate Vela X-1, which is a high-mass X-ray binary (HMXB) comprising a massive companion star and a neutron star Rawls et al. (2011) with observational values of radius $\mathfrak{R}=9.56_{-0.08}^{+0.08}$ km and mass $\mathcal{M}=1.77\pm 0.08~{}\mathcal{M}_{\odot}$. Sako et al. proposed an initial design for the Vela X-1’s global X-ray emission line spectrum Sako et al. (2000). Since its discovery, it has received a great deal of attention among the compact objects in this field of study. Here we consider the mass and radius of this compact object to develop a new physically stable and realistic stellar model in this modified gravity by investigating several stellar properties.
@@
This paper is organized as follows: In Section II, $f(R,T)$ gravity is defined mathematically in terms of isotropic matter distributions. A conceivable $f(R,\,T)$ gravity model is displayed in Section III. Section IV deals with boundary conditions; right here, we compare the interior and Schwarzschild’s exterior metrics to determine the values of the unknown constant for the selected values of our model parameters. Section V executes graphical analysis to investigate various physical properties and viability of compact stars. The last section is devoted to the conclusions. Throughout the investigation, we have mostly utilized a negative metric signature (+,,,)(+,-,-,-)( + , - , - , - ), and followed the gravitational or geometrized unit system ($c=1=G$) throughout our manuscript except for the calculations in Table 2.
@@ II Interior Spacetime and Basic field Equations
Now, the variation of the Ricci scalar provides the following expression:
@@
and the variation of the action term (1) can be written as:
@@
$f_{R}R_{\mu\nu}-\frac{1}{2}f(R,T)g_{\mu\nu}+(g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu})f_{R}$ $=$ $8\pi T_{\mu\nu}-f_{T}T_{\mu\nu}-f_{T}\Theta_{\mu\nu}$ (6)
@@
$\nabla^{\mu}T_{\mu\nu}$ $=$ $\frac{f_{T}}{8\pi-f_{T}}\left[(T_{\mu\nu}+\Theta_{\mu\nu})\nabla^{\mu}\ln f_{T}+\nabla^{\mu}\Theta_{\mu\nu}-\frac{1}{2}g_{\mu\nu}\nabla^{\mu}T\right].$ (7)
We can infer from the above equation that there is a chance that the matter energy-momentum tensor will not be conserved under this modified theory, in contrast to GR. An additional force is produced as a result of the geometry-matter coupling Koivisto (2006). This is a common characteristic of the $f(R,T)$ theories. This matter non-conservation may involve quantum effects and particle creation in cosmological scenario. This physical interpretation makes the most practical sense. However, the possibility of active matter-creation processes in the universe remains a hypothesis without any direct observational or experimental evidence. For a detailed discussion of this aspect in the context of $f(R,T)$ gravity, see Ref. Lobato et al. (2019). This non-conservation may occur when extending flat space expressions to curved space, resulting in a non-trivial conservation law. It is also possible to construct a conservative $f(R,T)$ gravity by making specific choices in the $f(T)$ counterpart, as illustrated in Ref. dos Santos Jr et al. (2019); Pretel et al. (2021); Carvalho et al. (2020). From Eqn.(7), we can easily check that the covariant derivative of the stress-energy tensor i.e. $\nabla^{\mu}T_{\mu\nu}\neq 0$ if $f_{T}(R,T)\neq 0.$ For this present model, $f_{T}(R,T)=2\chi$ and hence $\nabla^{\mu}T_{\mu\nu}\neq 0$. Hence, the system will not be conserved for our chosen $f(R,\,T)$ model. Moreover, when $f(R,T)=f(R)$, from eqn. (6) we derive the $f(R)$ gravity field equations.
@@
where $p$ denotes isotropic pressure and $\rho$ denotes matter density under modified gravity. The fluid velocity, $u^{\mu}$, fulfils the following equations: $u^{\mu}u_{\mu}=1$ and $u^{\mu}\nabla_{\nu}u_{\mu}=0$. As proposed by Harko et al. Harko et al. (2011), we have taken the matter Lagrangian that can be taken as $\mathcal{L}_{m}=-p$ and the expression of $\Theta_{\mu\nu}=-2T_{\mu\nu}-pg_{\mu\nu}.$
@@
$G^{0}_{0}:8\pi\rho^{\text{E}}$ $=$ $\frac{\lambda^{\prime}}{r}e^{-\lambda}+\frac{1}{r^{2}}(1-e^{-\lambda}),$ (12)
$G^{1}_{1}:8\pi p^{\text{E}}$ $=$ $\frac{1}{r^{2}}(e^{-\lambda}-1)+\frac{\nu^{\prime}}{r}e^{-\lambda},$ (13)
$G^{2}_{2}=G^{3}_{3}:8\pi p^{\text{E}}$ $=$ $\frac{1}{4}e^{-\lambda}\left[2\nu^{\prime\prime}+\nu^{\prime 2}-\lambda^{\prime}\nu^{\prime}+\frac{2}{r}(\nu^{\prime}-\lambda^{\prime})\right].$ (14)
Here $\rho^{\text{E}}$ and $p^{\text{E}}$ are respectively the density and pressure in Einstein Gravity and
$\rho^{\text{E}}$ $=$ $\rho+\frac{\chi}{8\pi}(3\rho-p),$ (15)
$p^{\text{E}}$ $=$ $p-\frac{\chi}{8\pi}(\rho-3p),$ (16)
the prime () indicates differentiation with respect to ‘r’. Using Eqs. (12)-(14), we get,
@@ III Exact Solution of our proposed Model for isotropic Stars
Utilizing the following metric potentials, one can compute the Einstein density and pressure as:
$\rho^{\text{E}}$ $=$ $\frac{3B\bigg{\{}(3+Br^{2})(1+4Br^{2})^{3/2}+C(3+9Br^{2})\bigg{\}}}{16\pi(1+Br^{2})^{2}(1+4Br^{2})^{3/2}},$ (20)
$p^{\text{E}}$ $=$ $-\frac{3B\bigg{\{}C+7BCr^{2}+3(-1+Br^{2})\sqrt{1+4Br^{2}}\bigg{\}}}{16\pi(1+Br^{2})^{2}\sqrt{1+4Br^{2}}},$ (21)
The physical parameters, like as Einstein pressure and density, must be obtained before so that we can investigate the full structure of stellar models.
By utilizing the $p^{\text{E}}$ and $\rho^{\text{E}}$ expressions from equations (15) and (16), we can derive the matter density and pressure expressions $\rho,\,p$ in modified gravity as,
@@ IV Exterior line element and matching conditions
$A$ $=$ $\sqrt{\frac{-2\mathcal{M}+\mathfrak{R}}{\mathfrak{R}(1+B\mathfrak{R}^{2})^{3}}}~{}~{}[{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\text{Considering only positive signed expression}}],$ (29)
$C$ $=$ $\frac{(1+4B\mathfrak{R}^{2})^{3/2}\Big{(}3\chi+6\pi-2B(\chi+3\pi)\mathfrak{R}^{2}\Big{)}}{2\pi+2B(3\chi+11\pi)\mathfrak{R}^{2}+7B^{2}(3\chi+8\pi)\mathfrak{R}^{4}},$ (30)
By fixing a preferred value of $B$ ($=0.001~{}\rm{km}^{-2}$), the values of $A$ and $C$ have been listed in Table 1 for various values of $\chi$, determined using the approximate mass and radius of the compact stellar candidate Vela X-1. From the expression of A in (29), we see that it is a function of $M,R$, and $B$. So, ultimately, A becomes a constant for this particular stellar model Vela X-1 after fixing a particular value of $B$. Thus $A$ is independent of $\chi$ throughout our model. On the contrary, the values of $C$ increase as $\chi$ increases.